OpenStudy (anonymous):

Hi! I would like to know how to solve: log 3=.5, then log 1/9=? If anyone could help, that be would be splendid! (:

OpenStudy (anonymous):

log 3=0.5 change that into exponential form

OpenStudy (anonymous):

\[Log_{10}3=0.5\]

OpenStudy (anonymous):

I don't know how to do that?

OpenStudy (anonymous):

\[Log_{b}a=c\]\[b^c=a\]

OpenStudy (anonymous):

Using the formula above \[10^{0.5}=3\]

OpenStudy (anonymous):

I have no clue! :/

OpenStudy (anonymous):

Tell me anything you know about this; anything

OpenStudy (anonymous):

Honestly I don't know anything, other than that this problem is for a math entrance exam for RU, and I haven't had a math class in well over 2 years...but the problem actually reads log a 3= 0.5, then log a 1/9=?

OpenStudy (anonymous):

No problem, we will start from beginning then

OpenStudy (anonymous):

Thanks! (:

OpenStudy (rosey):

u should be a fan of me <3 :)

OpenStudy (anonymous):

I will be! <3 (;

OpenStudy (anonymous):

I will start with my own example

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

\[3^x=9\] Based on your intuition what is x

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

right, how did you figure that out?

OpenStudy (anonymous):

Because 3 to the second power equals 9..

OpenStudy (rosey):

wooo thank u :)

OpenStudy (anonymous):

no problem(:

OpenStudy (anonymous):

Since it was obivous that 3^2=9 , you were able to figure that out What if it is 3^x=18

OpenStudy (anonymous):

In cases that are not obvious we use log to solve

OpenStudy (anonymous):

I don't know..I was thinking it might be 3^3, but that equals 27...so I'm absolutely clueless..

OpenStudy (anonymous):

Yes you can't by trial and error

OpenStudy (anonymous):

So we use log to solve

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

So our example is 3^x=18 we want to know x To do so we convert to log form

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

This is how it converts \[a^b=c---> Log_a c=b\]

OpenStudy (anonymous):

Trying converting our example 3^x=18 using form above

OpenStudy (anonymous):

so, do I switch the value placements of b with c then?

OpenStudy (anonymous):

3^18=x?

OpenStudy (anonymous):

no, \[Log_3 18=x\]

OpenStudy (anonymous):

Okay, I think I got it.

OpenStudy (anonymous):

Now, you can do reverse process as well \[Log_ab=c ---> a^c=b\]

OpenStudy (anonymous):

Using form above, try converting log 3=.5 to exponential form

OpenStudy (anonymous):

\[Log_{10} 3=.5\]

OpenStudy (anonymous):

where did the 10 come from?

OpenStudy (anonymous):

or did you just add that?

OpenStudy (anonymous):

When nothing is mentioned, we assumed it is ten.

OpenStudy (anonymous):

okay. 10^.5=3?

OpenStudy (anonymous):

wait,are you sure you question is right?

OpenStudy (anonymous):

you mean my answer?...if that's what you mean, then I have no clue. I'm sorry.

OpenStudy (anonymous):

forget the first one, solve the second one

OpenStudy (anonymous):

Is what you typed to me before, 3^x=18, converts into Log3 18=x?

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

okay...so you want me to try to solve log10 3=.5?

OpenStudy (anonymous):

you can't because it is like solving 7=2 does not make sense

OpenStudy (anonymous):

So, it's an unsolveable question?

OpenStudy (anonymous):

yes, makes no sense

OpenStudy (anonymous):

oh, I though I could just convert it the same way i converted it for the first formula, but i guess not/..but why?

OpenStudy (anonymous):

because there is no x for us to solve

OpenStudy (anonymous):

I see!

OpenStudy (anonymous):

second one press Log 1/9 in you calculuator

OpenStudy (anonymous):

I'm not sure if I input it in correctly, but I got -.954?

OpenStudy (anonymous):

Oh! And I have a question...so the x would also be considered a?

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

so it would be -2.2?

OpenStudy (anonymous):

no, I mean like in the problem 3^x=18->Log3 18=x, which can be substituted for a?

OpenStudy (anonymous):

oops that is not right let me do it correctly

OpenStudy (anonymous):

\[\log(\frac{1}{9})=\log(3^{-2})=-2\log(3)=-2\times \frac{1}{2}=-1\]

myininaya (myininaya):

is it \[\log_x(3)=\frac{1}{2}\]

OpenStudy (anonymous):

that is the solution to Hi! I would like to know how to solve: log 3=.5, then log 1/9=? If anyone could help, that be would be splendid!

OpenStudy (anonymous):

gimmick is to write \[\frac{1}{9}\] as a power of 3 and then use laws of logarithms. that is all

OpenStudy (anonymous):

satellite73: where did the -2 come from? Thanks! (:

OpenStudy (anonymous):

don't know what this other business is about, maybe correct but not necessary

myininaya (myininaya):

he was using properites of log

OpenStudy (anonymous):

\[3^{-2}=\frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (anonymous):

gotcha!

OpenStudy (anonymous):

that is where the -2 came from, recognizing that \[\frac{1}{9}=3^{-2}\]

myininaya (myininaya):

so did you get this question from an online thingy?

OpenStudy (anonymous):

Yeah...I believe I understand this problem now. Thanks! I appreciate it!! ...um yeah from the RUtgers website for a math placement exam... :/

OpenStudy (anonymous):

typical math teacher question. if \[\log(2)=.6\] and \[\log(5)=.8\] then \[\log(20)=?\] and you are supposed to say \[20=2^2\times 5\] so \[\log(20)=\log(2^2\times 5)=2\log(2)+\log(5)=2\times .6+.8\] etc

OpenStudy (anonymous):

I didn't have quite the best pre-calc teacher in high school, but I guess I could blame it on my own lack of initiative to do well, so I'm re-learning the basics again! :D

OpenStudy (anonymous):

or \[\log(2)=.6\] then \[\log(\frac{1}{16})=?\] and you are supposed to come up with \[2^{-4}=\frac{1}{16}\] so \[\log(\frac{1}{16})=\log(2^{-4})=-4\log(2)=-4\times .6\] andso on

myininaya (myininaya):

im just trying to figure out what base they meant because log base 10 of 3 does not =5

OpenStudy (anonymous):

sorry...I have no clue what that was about...I need to brush up on my logarithms and precalculus mathematical functions and all that... but thanks again for helping me solve the problem! (: The answer is -1

myininaya (myininaya):

i mean we can solve for the base

OpenStudy (anonymous):

it is not a known base. doesn't matter what base is

myininaya (myininaya):

but what they have isn't true

myininaya (myininaya):

it bothers me lol i mean why would they even give that part

OpenStudy (anonymous):

it says "IF \[\log(3)=\frac{1}{2}\]

OpenStudy (anonymous):

must be some base for which that is true. only wanted to see if you knew to write \[\frac{1}{9}=3^{-2}\] and use property of log

myininaya (myininaya):

yes thats untrue they should said log base x of 3 =1/2 that way it wouldn't be false

OpenStudy (anonymous):

i guess they should write \[\log_b(3)=\frac{1}{2}\] if only to make you happy

myininaya (myininaya):

yes! but the part is unnecessary at least it isn't untrue

OpenStudy (anonymous):

in fact the base is 9

OpenStudy (anonymous):

The problem states: given loga 3=.5, then loga (1/9)=? and the answer is -1

myininaya (myininaya):

lol

OpenStudy (anonymous):

yes that is the question and that is the answer by what i wrote above.

OpenStudy (anonymous):

whether we knew the base was 9 or not we still get answer \[\log_9(3)=\frac{1}{2}\] and \[\log_9(\frac{1}{9})=-1\]

myininaya (myininaya):

well that wasn't the question i seen but still you don't need the first part for the second part

OpenStudy (anonymous):

Okie dokie! thanks guys! (:

myininaya (myininaya):

\[\log_9(1)-\log_9(9)=0-1=-1\] since\[9^0=1and 9^1=9\]

OpenStudy (anonymous):

imran, you are supposed to solve for the base in the first equation I think.